### v2

# Dynamic Rank Adjustment for Accurate and Efficient Neural Network Training

Abstract
Low-rank training methods reduce the number of trainable parameters by re-parameterizing the weights with matrix decompositions (e.g., singular value decomposition). However, enforcing a fixed low-rank structure caps the rank of the weight matrices and can hinder the model’s ability to learn complex patterns. Furthermore, the effective rank of the model’s weights tends to decline during training, and this drop is accelerated when the model is reparameterized into a low-rank structure. In this study, we argue that strategically interleaving full-rank training epochs within low-rank training epochs can effectively restore the rank of the model’s weights. Based on our findings, we propose a general dynamic-rank training framework that is readily applicable to a wide range of neural-network tasks. We first describe how to adjust the rank of weight matrix to alleviate the inevitable rank collapse that arises during training, and then present extensive empirical results that validate our claims and demonstrate the efficacy of the proposed framework. Our empirical study shows that the proposed method achieves almost the same computational cost as SVD-based low-rank training while achieving a comparable accuracy to full-rank training across various benchmarks.
1 Introduction
Low-rank reparameterization methods have been actively studied for efficient training of large neural networks. Low-rank training strategies typically reduce the number of trainable parameters by applying matrix decompositions to a model’s weight matrices. While this approach lowers training cost, it permanently caps the maximum attainable rank of those matrices, thereby limiting the model’s ability to learn complex patterns. Moreover, recent studies report that the effective rank of the weights tends to decline during training [32, 11]. Therefore, to train large neural networks efficiently while preserving their learning capacity, it is crucial to address the decline in weight-matrix rank that occurs during low-rank training.
Singular value decomposition (SVD)–based method [14] is among the most widely used re-parameterization techniques. Other tensor factorizations, such as Tucker [18] and CP [20] decompositions, have likewise been adopted to enable low-rank training. In federated-learning settings, FLANC [23] and FedPara [12] re-parameterize models into low-rank forms. Although these approaches reduce the number of trainable parameters, they often ignore the progressive decline in the weights’ effective rank. Some regularization methods, such as soft orthogonal regularizer [32] and its variants [1, 17], focus on the rank decline issue. However, they incur higher computational cost and memory overhead, making them less practical for training large neural networks.
This study explores how to mitigate the decline in the effective rank of model weights in low-rank training. Our key finding is that interleaving a few full-rank epochs within low-rank training effectively preserves the model’s effective rank. Specifically, we analyze how run-time adjustments to the rank of model weights affect training and present a practical strategy for raising or lowering rank during low-rank training to maximize the model’s effective rank. Our theoretical analysis and empirical study demonstrate that dynamic rank adjustment matches the accuracy of full-rank training while retaining the system efficiency of low-rank methods.
Another key finding is that the effectiveness of rank adjustment is closely coupled with the learning rate. Our study shows that full-rank epochs should be scheduled according to the noise scale induced by the learning rate. In general, the learning rate tends to be initialized to a large value and then progressively decayed during training. In this study, we investigate how the learning rate influences the gap between low-rank and full-rank updates and, based on this analysis, develop a general rank scheduling framework that maximizes the benefits of interleaving full-rank epochs within low-rank training epochs. Table 1 shows a feature-wise comparison among various low-rank training methods.
To evaluate the performance of our proposed dynamic-rank training framework, we conduct an extensive empirical study on widely used computer vision and natural language processing benchmarks. We further benchmark our approach against state-of-the-art (SOTA) low-rank training techniques and regularization methods designed to restore effective rank. Throughout all the experiments, we find that our dynamic-rank training method achieves comparable accuracies to the full-rank training while significantly reducing the computational cost like conventional low-rank training methods. In addition, our experiments demonstrate that the proposed method integrates seamlessly with low-rank training when paired with soft-orthogonality (SO) regularization, confirming that the two techniques are complementary. we therefore conclude that the dynamic rank adjustment is a practical and easy-to-use training strategy that effectively boosts model accuracy in low-rank training.
2 Background
Method PE $R$ PC NA
Low-rank (SVD, etc.)
GKPD
PELA
TKD-CPD
Soft Orthogonality (SO)
SRIP
Dynamic-rank
Table 1: A feature-wise comparison between dynamic-rank training and other low-rank training methods. PE: parameter-efficient, $R$: rank recovery, PC: pre-training compatible, and NA: noise-scale aware.
Low-rank Re-parameterization – Low-rank reparameterization is a model approximation technique that is popularly used in large neural network training. Given a model weight matrix $\mathbf{W}$, a matrix decomposition method is applied to $\mathbf{W}$ before training begins. For example, if SVD is used, $\mathbf{W}\in\mathbb{R}^{m\times n}$ is decomposed to two smaller matrices $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ such that $\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$, where $k<m$ and $k<n$. As $k$ decreases, the total number of model parameters is reduced, thereby lowering the computational cost of training. However, the reconstructed weight matrix can have up to $k$ ranks, resulting in the limited learning capability. Low-rank Adaptation (LoRA) [10] is one of the most popular applications of re-parameterization technique, particularly in the context of fine-tuning. The importance of singular values under LoRA’s context is deeply analyzed in  [16].
Learning Rate and Noise Scale – In previous works, learning rate is known to play a key role in determining generalization performance of machine learning models. It has been theoretically shown that the noise scale $g$ in gradient approximation is determined by the learning rate and batch size, as follows [27]: $g\approx\eta N/B$, where $\eta$ is learning rate, $N$ is the dataset size, and $B$ is the mini-batch size. This relation links learning rate schedules to the notion of noise scale. In early training, a large noise scale helps the model explore and shape the decision boundary [22, 21], consistent with the common practice of decaying the learning rate. In this study, we investigate the relationship between the noise scale and the effective rank of the model’s weights.
Figure 1: Comparison of the layer-wise singular-value spectral ratio ($\lambda$) across different model ranks during ResNet20 training on CIFAR-10. The left plot shows layer-wise $\lambda$ curves for the full-rank model training while the right plot shows the $\lambda$ curves for the reparameterized model training. All convolution layers were reparameterized to have half rank. At the end of the training, many layers in the reparameterized model have large $\lambda$ values, meaning the model has converges to a low rank space.
3 Related Work
Low-rank Reparameterization Methods – Several studies have explored the use of low-rank re-parameterization techniques including singular value decomposition, Tucker decomposition, and CP decomposition methods [14, 18, 20]. Recently, GKPD proposed to re-parameterize the model weights using Kronecker-product decomposition and demonstrated promising performance [7]. PELA applies low-rank training to the pre-training phase to reduce the computational cost of training [6]. FLANC uses a customized tensor decomposition method to enable direct model aggregations across clients in the context of federated learning [23]. FedPara combines decomposition methods and the Hadamard product to enhance the representation capacity of re-parameterized models [12]. TKD-CPD jointly utilize Tucker and CP decompositions [25]. While these methods effectively reduce the number of trainable parameters and thus the computational and communication costs, they do not consider the inherent rank decline issue.
Rank Recovery Methods – Some regularization methods have been proposed to address the issue of inherent rank decline. Soft orthogonality (SO) is the basic regularization method which minimizes the difference between Gram matrix of the weight matrix and the identity matrix [32]. Double soft orthogonality (DSO) is a variant of SO which considers the regularization with overcomplete and undercomplete Gram matrices [1]. Spectral Restricted Isometry Property (SRIP) is a variant of RIP [2] that minimizes the spectral norm of the difference between the Gram matrix of a weight matrix and the identity matrix. Another study propose a sine-activated low-rank training strategy that is also designed to restore the model weight’s rank [15]. These methods commonly recover the effective rank of model weights, however, the rank cannot exceed the hard limit imposed by the low-rank reparameterization. In this study, we focus on how to overcome this limitation by adjusting the model rank at run-time.
3.1 Motivation
Recently, it has been theoretically shown that the rank of gradients tends to decrease as training progresses [11, 32]. Consequently, given a fixed training dataset and a large number of repeated training steps, the model weights are also expected to exhibit a similar reduction in rank.
As the singular value spectrum of model weights becomes more skewed, only a few singular vectors capture most of the information, while those associated with smaller singular values contribute little to the model’s representational capacity. We empirically verify this tendency by analyzing the effective rank of neural networks throughout training. To quantify how rapidly the model loses rank, we define layer-wise singular value spectral ratio, $\lambda^{l}$, as follows.
$\lambda^{l}=\frac{\sigma^{l}_{\max}}{\sigma^{l}_{\min}},\quad l\in[L],$ (1)
where $[L]$ denotes the set of all network layers, and $\sigma^{l}_{\max}$ and $\sigma^{l}_{\min}$ are the maximum and minimum singular values of layer $l$, respectively. Assuming that $\sigma^{l}_{\max}$ remains reasonably small, an increase in $\lambda^{l}$ indicates that $\sigma^{l}_{\min}$ is approaching zero, meaning the model is losing its rank.
Figure 1 shows how $\lambda^{l}$ changes during the training of ResNet20 on CIFAR-10. On the left-side (full-rank model training), only three layers show noticeable rank reduction while the effective ranks of all other layers remain quite stable throughout the training. On the right side (reparameterized to half rank at all convolution layers), most layers exhibit much higher $\lambda$ values compared to the full-rank training, resulting in a final model with low ranks in most layers. The full rank training achieves a validation accuracy of $92.23\%$, while the low-rank reparameterized training yields $91.08\%$. Based on this observation, we conclude that low-rank model reparameterization can significantly reduce the rank of model weights, thereby harming the model’s representational capacity.
Based on this empirical study, we can derive one critical insight as follows.
As the model is reparameterized to a lower rank,
the rank of its weights tends to decrease more rapidly.
Therefore, running full-rank training during the middle of training is expected to mitigate the rapid decline in the effective rank of model weights. This insight motivates the design of a general model rank adjustment framework, which we describe in the following section.
4 Dynamic-rank Training Framework
In this section, we propose a general dynamic-rank training framework that temporarily increases the rank of the model’s weight matrices at selected points during training. We first formalize rank inflation and deflation, then discuss how to schedule rank adjustments. In particular, we provide two key insights for placing full-rank epochs within low-rank training to maximize rank restoration. Figure 2 illustrates the proposed framework.
Figure 2: A schematic illustration of dynamic-rank training framework.
4.1 Model Rank Adjustment
We define two directions of model rank adjustment, inflation and deflation, as follows. For simplicity, we use the notation of SVD-style low-rank reparameterization, but it can be easily replaced with other methods such as Tucker or CP decompositions.
Model Inflation – Low-rank training can be implemented in two different forms, vanilla low-rank reparameterization and low-rank adaptation, LoRA [10]. Without loss in generality, we define model inflation process as follows. Given low-rank model weights $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ and base parameter $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$, the model is inflated such that $\mathbf{W}\leftarrow\mathbf{W}_{0}+\mathbf{A}\mathbf{B}^{\top}$. Note that we assume $k<r\leq n$. If the low-rank model follows the standard low-rank reparameterization, the initial weight matrix $\mathbf{W}_{0}$ is set to the zero matrix, i.e., $\mathbf{0}_{m\times n}$. Consequently, the maximum available rank is increased from $k$ to $n$. Please see Appendix for the case of convolutional layers.
Model Deflation – Given a model weight $\mathbf{W}$, the model is deflated by attaching a low-rank adaptor path next to the original weight such that $\mathbf{W}_{f}+\mathbf{A}\mathbf{B}^{\top}\leftarrow\mathbf{W}$, where $\mathbf{W}_{f}\in\mathbb{R}^{m\times n}$ is the given model weight matrix and $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are low-rank model weights. The provided weight matrix is frozen as $\mathbf{W}_{f}$ and $\mathbf{A}$ and $\mathbf{B}$ are trained instead. One can initialize $\mathbf{A}$ and $\mathbf{B}$ using either random distributions or zero matrices. Note that this naturally resembles LoRA when SVD is used as the reparameterization method.
These model rank adjustment procedures incur additional computations during training, however, when the epoch budget $E$ is sufficiently large, the extra computational cost typically becomes negligible. We now turn our attention to when the model rank should be inflated or deflated during training.
4.2 Model Rank Scheduling
Let the total training budget be $E$ epochs. We begin training in a low-rank form and increase the rank of model weights at epoch $I$. Then, at epoch $D$, we reduce the rank again. That is, training is performed in the full-rank space for DID-I epochs, and in the low-rank space for the remaining E(DI)E-(D-I) epochs. Under this setting, the key question is how to find good $I$ and $D$. Here, we present two empirical insights that motivate practical principles for selecting $I$ and $D$ to maximize rank recovery.
Figure 3: CIFAR-10 (ResNet20) benchmark with various dynamic-rank training schedules. Inflate and Deflate indicate the epoch where the model rank is increased and decreased, respectively. The best accuracy is achieved when the full-rank epochs are located in both high-noise and low-noise regimes. By contrast, restricting full-rank epochs to only one of these regimes leaves the training loss insufficiently reduced.
Full-rank training should be performed in both high-noise and low-noise regimes – We propose performing full-rank training both before and after the learning rate decay. This ensures that all ranks are sufficiently exposed to different learning rates during training. After deflating the model, only up to $k$ ranks are updated in the adaptor and the remaining nkn-k ranks remain unchanged in the frozen unless the model is inflated again. Thus, if the model is deflated before the learning rate decay and not inflated again in the rest of the training, the nkn-k ranks will not be trained under the low-noise regime. Those nkn-k ranks may lead to underfitting, resulting in lower model accuracy.
In contrast, if full-rank training is performed only in the low-noise regime, the model is likely to suffer from poor generalization performance. It has been shown that training in a high-noise regime during the early phase improves generalization performance [22, 8, 21]. Therefore, if the model is inflated for the first time in the low-noise regime, the nkn-k ranks will be trained only with a small learning rate, potentially degrading the model’s generalization performance. Moreover, scheduling the full-rank epochs later reduces the training iterations for the adapter’s $k$ ranks, potentially causing underfitting.
Figure 3 shows simple empirical results which validates our argument. We evaluated CIFAR-10 with ResNet-20 under three schedules that insert the full-rank training epochs at different points. The model is reparameterized with SVD, reducing its rank by half. Because the learning rate decays after epoch 100, the first and third settings correspond to full-rank epochs only in the high-noise and low-noise regimes, respectively. As expected, the best accuracy is achieved when full-rank epochs are located in both noise regimes. Note that the curves spike during deflation, as a randomly initialized low-rank adaptor is attached and begins training. The yellow and orange curves clearly show both underfitting and low accuracy. Based on this analysis, we conclude that full-rank training should be applied in both the high-noise and low-noise regimes to sufficiently train the model.
Full-rank training should be at the end of the high-noise regime and the early low-noise regime – We begin by analyzing the rank boundaries of the reconstructed model after deflation, as below.
Setting II $D$ Acc. $\lambda^{0}$ $\lambda^{10}$ $\lambda^{18}$
None - - $91.14\%$ $\infty$ $7.58$ $\infty$
Early 0 50 $91.26\%$ $59.32$ $6.02$ $76.47$
Middle 25 75 $91.82\%$ $39.29$ $4.27$ $74.18$
Late 50 100 $91.87\%$ $37.51$ $4.19$ $70.36$
Table 2: Layer-wise comparison of singular-value spectral ratio ($\lambda$) across different rank-adjustment schedules. The $I$ and $D$ indicate rank inflation and deflation, respectively. ResNet20 was trained on CIFAR-10 for 150 epochs, with the learning rate reduced by a factor of 10 at epochs 100 and 130. Having full-rank epochs at the end of the high-noise regime yields the best results.
Proposition 1. Suppose model weight matrices: $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$, $\mathbf{A}\in\mathbb{R}^{m\times k}$, and $\mathbf{B}\in\mathbb{R}^{n\times k}$ and a re-constructed matrix $\mathbf{W}=\mathbf{W}_{0}+\mathbf{A}\mathbf{B}^{\top}$. Assuming that the effective rank of $\mathbf{W}_{0}$ is $r$, the rank of the reconstructed matrix $\mathbf{W}$ is bounded between rkr-k and $r+k$.
Proof.
First, re-write each matrix as a summation of rank-1 matrices such that $\mathbf{W}_{0}=\sum_{j=1}^{r}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}$ and $\mathbf{A}\mathbf{B}^{\top}=\sum_{j=1}^{k}\mathbf{a}_{j}\mathbf{b}_{j}^{\top}$. Consequently, $\mathbf{W}=\sum_{j=1}^{r}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}+\sum_{j=1}^{k}\mathbf{a}_{j}\mathbf{b}_{j}^{\top}$. The minimum rank of $\mathbf{W}$ is rkr-k if the $k$ rank-1 matrices in $\mathbf{A}\mathbf{B}^{\top}$ eliminate $k$ rank-1 components of $\mathbf{W}_{0}$. On the other hand, the maximum rank of $\mathbf{W}$ is $r+k$ if the $k$ rank-1 matrices in $\mathbf{A}\mathbf{B}^{\top}$ are orthogonal to all $k$ components of $\mathbf{W}_{0}$. ∎
This analysis suggests that the rank of the reconstructed model weights is closely tied to the rank of $\mathbf{W}_{0}$, the frozen weights. In particular, performing full-rank training right before deflating the model maximizes $r$, enabling the model to maintain a high effective rank. The full-rank training budget, DID-I, may be smaller than the number of epochs before the first learning rate decay. In such cases, training will naturally begin with a low-rank reparameterized model, which will then be inflated before the learning rate decay.
To connect theory and practice, we compare weight ranks under different rank schedules in Table 2. We perform CIFAR-10 (ResNet-20) training with the learning rate decaying at epoch 100. The rightmost three columns show $\lambda$ for layers 0, 10, and 18 (near the input, middle, and output). In case 1 (conventional low-rank), $\lambda_{0}$ and $\lambda_{18}$ are very large, $\infty$ denotes values $10^{6}$, indicating widely spread spectra and effective rank loss. As full-rank training is applied later, the spectra narrow (smaller $\lambda$ values). These results demonstrate that the full-rank budget should be placed toward the end of the high-noise regime.
Now, let us discuss the impact of full-rank training in the low-noise regime. Once the learning rate decays, the rank of the model weights does not change dramatically due to the reduced magnitude of updates. Instead, what matters more is the extent to which low-rank training pushes the model away from the full-rank convergence point. The smaller the learning rate, the less significant the difference between low-rank and full-rank training becomes.
Proposition 2. As the learning rate decays, the gap between the full-rank model update and the low-rank model update is likely to decrease.
Proof.
The gradient of low-rank model weights, $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ can be written as $\nabla f(\mathbf{A}\mathbf{B}^{\top})\mathbf{B}\in\mathbb{R}^{m\times k}$ and $\nabla f(\mathbf{A}\mathbf{B}^{\top})^{\top}\mathbf{A}\in\mathbb{R}^{n\times k}$, respectively. Therefore, the reconstructed model can be written as follows.
$\mathbf{W}_{t+1}$ $=\mathbf{A}_{t+1}\mathbf{B}_{t+1}^{\top}$
$=\left(\mathbf{A}_{t}-\eta\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\right)\left(\mathbf{B}_{t}-\eta\nabla f(\mathbf{W}_{t})^{\top}\mathbf{A}_{t}\right)^{\top}$
$=\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$
$\hskip 14.22636pt-\eta\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{B}_{t}^{\top}-\eta\mathbf{A}_{t}\mathbf{A}_{t}^{\top}\nabla f(\mathbf{W}_{t})$
$\hskip 14.22636pt+\eta^{2}\nabla f(\mathbf{W}_{t})\mathbf{B}_{t}\mathbf{A}_{t}^{\top}\nabla f(\mathbf{W}_{t}).$
Since $\mathbf{A}_{t}\mathbf{B}_{t}^{\top}=\mathbf{W}_{t}$, the remaining three terms on the right-hand side can be considered as the low-rank update. For simplicity, we shorten the terms as follows.
$\mathbf{X}$ $:=\nabla f(\mathbf{W})\mathbf{B}\mathbf{B}^{\top}$
$\mathbf{Y}$ $:=\mathbf{A}\mathbf{A}^{\top}\nabla f(\mathbf{W})$
$\mathbf{Z}$ $:=\nabla f(\mathbf{W})\mathbf{B}\mathbf{A}^{\top}\nabla f(\mathbf{W})$
Then, the difference between the full-rank update $\nabla f(\mathbf{W}_{t})$ and the low-rank update, $d$ is defined as follows.
$d_{t}:=\left\|\nabla f(\mathbf{W}_{t})-\left(-\mathbf{X}_{t}-\mathbf{Y}_{t}+\eta\mathbf{Z}_{t}\right)\right\|_{F}$ (2)
Based on triangle inequality, (2) is upper-bounded as follows.
$d_{t}$ $\leq\left\|\nabla f(\mathbf{W}_{t})\right\|_{F}$
$\hskip 14.22636pt+\left\|\mathbf{X}_{t}\right\|_{F}+\left\|\mathbf{Y}_{t}\right\|_{F}+\eta\left\|\mathbf{Z}_{t}\right\|_{F}.$ (3)
Then, $\|\mathbf{X}_{t}\|_{F}$ on the right-hand side can be further bounded as follows.
$\mathbf{X}_{t}$ $=\|\nabla f(\mathbf{W}_{t})\mathbf{B}\mathbf{B}^{\top}\|_{F}$
$\leq\|\nabla f(\mathbf{W}_{t})\|_{F}\cdot\|\mathbf{B}\mathbf{B}^{\top}\|_{2}$
$=\|\nabla f(\mathbf{W}_{t})\|_{F}\cdot\|\mathbf{B}\|_{2}^{2}$
The same procedure can be applied to the $\mathbf{Y}$ and $\mathbf{Z}$ terms, yielding
$d_{t}$ $\leq\|\nabla f(\mathbf{W}_{t})\|_{F}\left(1+\|\mathbf{A}_{t}\|_{2}^{2}+\|\mathbf{B}_{t}\|_{2}^{2}\right)$
$\hskip 14.22636pt+\eta\|\nabla f(\mathbf{W}_{t})\|_{F}\left(\|\nabla f(\mathbf{W}_{t})\|_{2}\cdot\|\mathbf{A}_{t}\|_{2}\cdot\|\mathbf{B}_{t}\|_{2}\right).$
According to the bound above, the maximum difference decreases as $\eta$ decays. ∎
Note that, for clarity, we omit $\mathbf{W}_{0}$ from the analysis. Incorporating $\mathbf{W}_{0}$ into each step of the proof to recover LoRA-style low-rank training is straightforward. The above analysis shows that full-rank training should be performed as early as possible in the low-noise regime to minimize $d_{t}$. In general, most popular learning rate schedules in neural network training monotonically decrease the learning rate over time. Therefore, performing full-rank training earlier tends to result in a smaller accumulated difference, $\sum_{t}d_{t}$, during the rest of the training epochs.
4.3 Unified Rank Adjustment Framework
Based on the two key insights discussed above, we build up a general model rank adjustment framework as shown in Figure 2. The training begins in a low-rank space. The model’s rank is adjusted twice throughout the training, increased in a high-noise regime and then decreased in a low-noise regime. In our empirical study, we observed reasonably good performance when $I$ is set to the middle of high-noise regime and $D$ to the middle of low-noise regime. For example, given $E=150$, if learning rate decays at epoch $100$, we recommend beginning with $I=50$ and $D=125$ and then fine-tune these values to minimize training time while maintaining model accuracy. We will include an ablation study to analyze how the choices of $I$ and $D$ affect model accuracy and computational cost in the following section.
5 Experiments
Method Rank Ratio $\rho$ CIFAR-10 (ResNet20) CIFAR-100 (WRN28-10) Tiny ImageNet (ResNet50)
Acc. Comp. Acc. Comp. Acc. Comp.
Full-Rank $1.00$ $92.15\pm 0.1\%$ $1.00$ $78.82\pm 0.1\%$ $1.00$ $60.32\pm 0.1\%$ $1.00$
SVD $0.50$ $91.19\pm 0.1\%$ $0.57$ $73.57\pm 0.1\%$ $0.56$ $54.90\pm 0.1\%$ $0.79$
$0.75$ $91.56\pm 0.1\%$ $0.84$ $76.90\pm 0.1\%$ $0.84$ $56.16\pm 0.1\%$ $0.92$
Dynamic Rank (SVD) $0.50$ $\mathbf{92.13}\pm 0.1\%$ $0.76$ $\mathbf{78.61}\pm 0.2\%$ $0.64$ $\mathbf{61.80}\pm 0.1\%$ $0.85$
Tucker $0.50$ $90.27\pm 0.1\%$ $0.41$ $68.77\pm 0.1\%$ $0.39$ $60.62\pm 0.1\%$ $0.69$
$0.75$ $91.50\pm 0.1\%$ $0.79$ $79.68\pm 0.2\%$ $0.77$ $61.07\pm 0.7\%$ $0.87$
Dynamic Rank (Tucker) $0.50$ $\mathbf{92.10}\pm 0.1\%$ $0.70$ $\mathbf{78.48}\pm 0.1\%$ $0.66$ $\mathbf{61.25}\pm 0.1\%$ $0.84$
CP $0.50$ $88.31\pm 0.1\%$ $0.31$ $65.55\pm 0.1\%$ $0.30$ $42.05\pm 0.1\%$ $0.65$
$0.75$ $88.90\pm 0.1\%$ $0.59$ $65.32\pm 0.1\%$ $0.57$ $43.12\pm 0.1\%$ $0.78$
Dynamic Rank (CP) $0.50$ $\mathbf{91.33}\pm 0.1\%$ $0.61$ $\mathbf{78.61}\pm 0.3\%$ $0.65$ $\mathbf{59.56}\pm 0.5\%$ $0.83$
TKD-CPD (Tucker + CP) 0.50 $87.99\pm 0.1\%$ 0.21 $58.90\pm 0.1\%$ $0.18$ $34.53\pm 0.1\%$ 0.60
0.75 $89.62\pm 0.1\%$ 0.35 $60.22\pm 0.1\%$ $0.31$ $34.84\pm 0.1\%$ 0.66
Dynamic Rank (TKD-CPD) 0.50 $\mathbf{91.24}\pm 0.1\%$ 0.55 $\mathbf{73.96}\pm 0.3\%$ 0.59 $\mathbf{60.06}\pm 0.5\%$ 0.80
Table 3: Validation accuracy comparison across various low-rank training methods. The detailed settings including the inflate (II) and deflate ($D$) epochs are provided in Appendix. Regardless of which low-rank decomposition technique is used, our method consistently achieves significantly higher accuracy while effectively reducing computational cost.
Experimental Settings – We evaluate dynamic rank training framework on three computer vision benchmarks: CIFAR-10 [19] (ResNet20), CIFAR-100 (Wide-ResNet28-10), and TinyImageNet [24] (ResNet50), and eight natural language processing (NLP) benchmarks in GLUE [29] (DeBERTaV3-base [9]). All experiments are conducted on a GPU server equipped with two NVIDIA RTX 4090 GPUs. Each experiment is repeated at least twice, and we report the average accuracy along with the standard deviation. Detailed hyperparameter settings are provided in the Appendix.
5.1 Comparative Study
Computer Vision Benchmarks – To evaluate the performance of our dynamic-rank training strategy, we compare it with several low-rank reparameterization methods: SVD [14], Tucker [18], and CP [20] decompositions. We also compare it with TKD-CPD [25], a recently proposed low-rank training method that jointly utilizes Tucker and CP decompositions. Table 3 shows the results. The Rank Ratio $\rho$ indicates the ratio of the reparameterized weight rank to the original model’s maximum rank. Comp. refers to the average proportion of trainable parameters during training. Across all benchmarks, low-rank training leads to noticeable drops in accuracy. While increasing $\rho$ to 0.75 narrows the gap, it still falls short of full-rank performance. In contrast, dynamic-rank training matches full-rank accuracy while maintaining a similar computational cost to the $\rho=0.75$ setting. While TKD-CPD dramatically reduces the computational cost, it yields the lowest model accuracy among all the low-rank methods. However, when dynamic-rank training is applied to TKD-CPD, its accuracy is restored to a level comparable to full-rank training. These results suggest that the dynamic-rank training framework is readily applicable to a variety of modern machine learning applications.
Method Comp. CoLA MNLI MRPC QNLI QQP RTE SST-2 SST-B Average
Full Fine-Tuning $1.000$ $0.6750$ $0.9018$ $0.8995$ $0.9397$ $0.9222$ $0.8375$ $0.9587$ $0.9097$ $0.8764$
LoRA $0.014$ $0.6333$ $0.8874$ $0.8799$ $0.9192$ $0.8922$ $0.7978$ $0.9461$ $0.9021$ $0.8618$
AdaLoRA $0.017$ $0.6599$ $0.8891$ $0.8431$ $0.9183$ $0.8891$ $0.8086$ $0.9477$ $0.8943$ $0.8564$
Dynamic Rank (SVD) $0.115$ $\mathbf{0.6752}$ $\mathbf{0.8959}$ $\mathbf{0.8886}$ $\mathbf{0.9264}$ $\mathbf{0.9062}$ $\mathbf{0.8267}$ $\mathbf{0.9481}$ $\mathbf{0.9087}$ $\mathbf{0.8714}$
Table 4: Test accuracy comparison with NLP model (DeBERTaV3-base) and GLUE benchmark.
Natural Language Processing Benchmarks – We validate dynamic-rank training framework by fine-tuning the DeBERTaV3-base on 8 datasets from GLUE benchmark [29]. In fine-tuning, there is no high-noise regime, as the learning rate is typically set to a small value and remains in the low-noise regime throughout training. Therefore, in these experiments, we place full-rank epochs at the beginning of fine-tuning.
Table 4 reports the results. We compare our method with LoRA [10], the most widely used low-rank fine-tuning method, and AdaLoRA [33], a representative variant. All low-rank methods greatly reduce compute versus full fine-tuning. LoRA is the most efficient but shows a noticeable accuracy drop. AdaLoRA mitigates the drop but still does not match the full fine-tuning accuracy. By contrast, dynamic-rank training achieves accuracy comparable to that of full fine-tuning. Although the computational cost is roughly $7$ times AdaLoRA’s cost but still far cheaper than full fine-tuning, making it a practical choice for NLP applications.
Dynamic-rank Training with Regularization – Our dynamic-rank training strategy is readily applicable to regularization methods designed to restore the reduced effective rank.
Table 5 shows CIFAR-10 benchmark results with three regularization methods, soft orthogonality (SO) [32], double soft orthogonality (DSO) [1], and SRIP+ [17]. These regularizers effectively improve the accuracy of SVD-based low-rank training. When combined with dynamic-rank training, they yield consistent accuracy gains across all regularizers. Therefore, the two are complementary, and their combination achieves high accuracy at low computational cost. However, these regularizers consume a large amount of memory space to compute Gram matrices, and the overall memory footprint may substantially increase as they are combined with dynamic-rank training.
Parameter Heatmap Comparison – Figure 4 shows a parameter value comparison between low-rank training and dynamic-rank training. We draw heatmaps with 4-D weight tensors of size $3\times 3\times 64\times 64$ obtained from the largest convolution layer in ResNet20. The tensors are reshaped to $192\times 192$ 2-D matrices. To show the difference more clearly, we reconstructed the tensor from the low-rank reparameterized weights using the 50 singular vectors corresponding to the smallest singular values. The low-rank model’s map exhibits blocky artifacts, whereas the dynamic-rank model’s map does not. The presence of blocky artifacts implies the existence of a null space in certain directions of the feature space, meaning that variations in the input along those directions do not affect the output. This result is well aligned with the model accuracy comparisons shown in other tables. Consequently, we conclude that the dynamic-rank training framework well restores the effective rank of the weight matrix, thereby improving model accuracy.
Setting Low-rank Dynamic Rank
Low-rank SVD (0.5) $91.14\pm 0.1\%$ $\textbf{92.09}\pm 0.1\%$
Low-rank SVD (0.5) + SO $91.61\pm 0.1\%$ $\textbf{92.35}\pm 0.2\%$
Low-rank SVD (0.5) + DSO $91.76\pm 0.4\%$ $\textbf{92.49}\pm 0.1\%$
Low-rank SVD (0.5) + SRIP+ $91.50\pm 0.0\%$ $\textbf{92.23}\pm 0.2\%$
Table 5: CIFAR-10 (ResNet20) accuracy comparison between with and without applying the proposed method to SOTA rank recovery methods.
5.2 Ablation Study
In our empirical study, we found that dynamic-rank training performs well when $D$ and $I$ are set to allocate a similar number of full-rank epochs to both the high-noise and low-noise regimes. We now conduct a simple ablation study to examine how the number of full-rank epochs affects model accuracy. Table 6 presents the results. SVD-based low-rank training is evaluated on the CIFAR-10 and CIFAR-100 benchmarks. For convenience, we define $\phi=(D-I)/E$ as the ratio of full-rank epochs to the total training budget. When $\phi=0.0$, it corresponds to conventional low-rank training; when $\phi=1.0$, it becomes full-rank training. On both benchmarks, accuracy starts to drop when $\phi$ goes below 0.5, particularly in the range $\phi\in[0.3,0.5]$. We therefore recommend starting with $\phi=0.5$ and adjusting downward as needed.
Figure 4: Parameter comparison between low-rank and dynamic-rank trainings. The heatmap shows the weights of the largest convolution layer in ResNet20 after training.
Setting Rank Comp. CIFAR-10 CIFAR-100
Ratio $\rho$
$\phi=0.1$ 0.25 0.36 $90.54\pm 0.3\%$ $76.29\pm 0.1\%$
$\phi=0.3$ 0.51 $91.43\pm 0.1\%$ $77.03\pm 0.4\%$
$\phi=0.5$ 0.64 $92.01\pm 0.3\%$ $78.61\pm 0.2\%$
$\phi=0.7$ 0.79 $92.08\pm 0.4\%$ $78.75\pm 0.1\%$
$\phi=0.9$ 0.93 $92.14\pm 0.2\%$ $78.63\pm 0.3\%$
$\phi=0.1$ 0.5 0.61 $91.28\pm 0.1\%$ $76.84\pm 0.1\%$
$\phi=0.3$ 0.69 $91.57\pm 0.1\%$ $77.51\pm 0.2\%$
$\phi=0.5$ 0.78 $92.11\pm 0.2\%$ $78.41\pm 0.3\%$
$\phi=0.7$ 0.87 $92.15\pm 0.2\%$ $78.39\pm 0.1\%$
$\phi=0.9$ 0.96 $92.11\pm 0.1\%$ $78.47\pm 0.3\%$
Table 6: Ablation study on how $\phi=(D-I)/E$ affects the performance of dynamic-rank training. Low-Rank SVD is used for all experiments. When $\phi=0.0$, it becomes the conventional low-rank training. In contrast, when $\phi=1.0$, it becomes full-rank training. The $I$ and $D$ are set to perform the same number of full-rank epochs in the high-noise and low-noise regimes.
5.3 Computational Cost Analysis
Here, we analyze the computational overhead of dynamic-rank training, which is shown in Comp. columns in Table 3 and 6. First, the full-rank training cost is calculated as
$T_{F}=E\cdot d,$ (4)
where $E$ is the total number of training epochs and $d$ is the model size. Then, the low-rank training cost is calculated as
$T_{SVD=0.5}=E\cdot d_{SVD=0.5},$ (5)
where the subscription $SVD=0.5$ indicates the model is reparameterized using SVD and the rank reduction ratio $\rho=0.5$. Finally, our dynamic-rank training cost is calculated as
$T_{DR}=(D-I)\cdot d+(E-(D-I))\cdot d_{SVD=0.5}.$ (6)
Thus, if $\rho$ is the same, $T_{DR}$ should be higher than $T_{SVD}$ and lower than $T_{F}$. For example, ResNet20 contains 272,762 trainable parameters and the reparameterized model with SVD ($\rho=0.5$) has 155,170 parameters. When $E=150$, $I=60$, and $D=135$, where $\phi=0.5$, $T_{DR}/T_{F}$ becomes $0.7844$, which is shown in Comp. column in Table 3. Our empirical study demonstrates that, by carefully tuning $I$ and $D$, a relatively small DID-I value provides comparable accuracy to the full-rank training. Small DID-I settings in dynamic-rank training makes its training cost more similar to that of the conventional low-rank training.
6 Conclusion
In this study, we propose a dynamic-rank training framework that restores the effective rank of model weights while preserving the efficiency of low-rank training. We also present two key insights for maximizing rank recovery by strategically interleaving full-rank epochs within low-rank training. Our extensive empirical study proves that scheduling full-rank epochs at the end of the high-noise regime and the beginning of the low-noise regime maximizes recovery of the model’s effective rank, thereby improving accuracy. The proposed dynamic-rank training scheme is general and readily applicable to any deep learning applications. We believe that harmonizing the proposed dynamic-rank training and other compute-efficient neural network training methods can be a promising future work.